OpenStudy (anonymous):

A softball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 5. h = height t = time in seconds Find the maximum height of the ball.

OpenStudy (anonymous):

take derivative & = 0

OpenStudy (anonymous):

h'=-32t+32=0 t=1 (sec) put the value in original equation & find h(1)

OpenStudy (tad1):

If you're not a calculus student, you can find the answer by finding the vertex of the parabola at (-b/2a, f(-b/2a)). where a = -16, b = 32, and c = 5.